### v1

@@ 5 Results
In conclusion, the experimental results provide strong evidence of the effectiveness of Curled-Dreamer in addressing various reinforcement learning tasks. The combination of DreamerV3’s predictive capabilities and the enhanced representation learning from CURL makes Curled-Dreamer a powerful and versatile algorithm for visual reinforcement learning.
@@ 6 Conclusion
The results indicate that the enhanced representation learning achieved through contrastive and reconstruction losses is key to Curled-Dreamer’s performance. These components enable the model to capture more informative and discriminative features from high-dimensional visual inputs, leading to better policy performance. The robust training procedure and well-tuned hyperparameters further contribute to the algorithm’s success, ensuring stability and efficiency in learning.
Furthermore, the addition of parameters, such as the use of two encoders and the matrix $W$, likely contributed to the improved results by allowing the model to capture more complex features and relationships within the data. This increase in model capacity, combined with effective training strategies, enhanced Curled-Dreamer’s ability to learn and generalize from visual inputs.

### v2

@@ 5 Results
In conclusion, the experimental results provide strong evidence of the effectiveness of Curled-Dreamer in addressing various reinforcement learning tasks. The combination of DreamerV3’s predictive capabilities and the enhanced representation learning from CURL makes Curled-Dreamer a new option for model-based reinforcement learning.
@@ 6 Conclusion
The results indicate that the representation learning achieved through contrastive and reconstruction losses is key to Curled-Dreamer’s performance. These components enable the model to capture more informative and discriminative features from high-dimensional visual inputs, leading to better policy performance. The robust training procedure and well-tuned hyperparameters further contribute to the algorithm’s success, ensuring stability and efficiency in learning.
Furthermore, the addition of parameters, such as the use of two encoders and the matrix $W$, likely contributed to the improved results by allowing the model to capture more complex features and relationships within the data. This increase in model capacity. enhanced Curled-Dreamer’s ability to learn and generalize from visual inputs.